# High-Performance Statistical Computing (HPSC): Challenges, Opportunities, and Future Directions

### v2

@@ Abstract
We recognize the emergence of a statistical computing community focused on working with large computing platforms and producing software and applications that exemplify high-performance statistical computing (HPSC). The statistical computing (SC) community develops software that is widely used across disciplines. However, it remains largely absent from the high-performance computing (HPC) landscape, particularly on platforms such as those featured on the www.top500.org or Green500 lists. Many disciplines already participate in HPC, mostly centered around simulation science, although data-focused efforts under the artificial intelligence (AI) label are gaining popularity. Bridging this gap requires both community adaptation and technical innovation to align statistical methods with modern HPC technologies. We can accelerate progress in fast and scalable statistical applications by building strong connections between the SC and HPC communities. We present a brief history of SC, a vision for how its strengths can contribute to statistical science in the HPC environment (such as HPSC), the challenges that remain, and the opportunities currently available, culminating in a possible roadmap toward a thriving HPSC community.
@@ 1 Introduction
Statistical computing (SC) is a foundational discipline that aims to merge statistical theory with computational techniques to turn data into actionable insights [Kennedy and Gentle, 2021]. At its core, it involves designing and implementing novel algorithms, simulations, and models to solve complex problems across various domains, including climate science, economics, and machine learning. Unlike general-purpose computing, statistical computing is grounded in data analysis, probabilistic reasoning, and statistical inference, drawing heavily on probability theory, statistics, and numerical methods [Givens and Hoeting, 2012]. As data continues to scale in both size and complexity, statistical computing has become indispensable, not only for managing large datasets but also for enabling predictive modeling and informed decision-making in real-time. Programming languages such as R [Ihaka and Gentleman, 1996], Julia [Bezanson et al., 2017], and Python [Van Rossum and Drake Jr, 1995] have played a critical role in this evolution, empowering researchers to rapidly post-process, analyze, and visualize data. Together, these tools have made SC a central engine of modern data science.
@@
To date, a lot more work within the SC community has been invested in adopting dataflow technologies, i.e., frameworks such as Apache Spark [Hariadi et al., 2020], Dask [Rocklin, 2015], or TensorFlow [Abadi et al., 2016], that represent computations as directed acyclic graphs of operations on data than in exploring message passing interface (MPI)+X approaches, i.e., a hybrid parallel programming paradigm that combines MPI [Gropp et al., 1999] for distributed-memory communication with an additional model “X” (such as OpenMP [Chandra, 2001] for multicore CPUs or CUDA for GPUs) for shared-memory or accelerator-level parallelism [Hori et al., 2021], despite the latter being better suited for many of the most computationally demanding tasks in the field. There are multiple factors behind this trend. First, dataflow architectures are often perceived as more accessible, thanks to their relatively simple parallel execution models, in contrast to MPI-based systems, which require explicit message passing and distributed memory management. Such frameworks abstract away much of this complexity. This helps statisticians use them even if they do not fully understand parallel programming. Second, the need for interactive computing has strongly influenced choices. Tools like R and Python are deeply tied to exploratory, iterative analysis, where users develop models, visualize outputs, and debug interactively. Moreover, dataflow systems integrate more naturally into cloud-based and batch-processing environments, making them a compelling choice for both industry and large-scale analytics. Third, most statisticians and data scientists primarily work with multicore processors rather than large HPC clusters. This has led to a strong emphasis on shared-memory parallelism (e.g., OpenMP, threading in R and Python) rather than distributed memory paradigms, such as MPI.
While dataflow technologies offer some distributed computing capabilities, they are often employed in a manner that still aligns with a multicore approach rather than fully leveraging the power of supercomputing. Even when statisticians engage with distributed computing, it is often through tools such as Dask [Rocklin, 2015], Ray [Moritz et al., 2018], or Spark, which offer a gentler learning curve and allow scaling across nodes without requiring low-level programming expertise. While these tools have clear advantages in usability and accessibility, we confirm that MPI+X remains significantly underutilized in SC. Supercomputing environments offer high efficiency, lower latency, and access to highly optimized numerical libraries. These advantages can substantially accelerate tasks such as large-scale simulations, Monte Carlo methods, and Bayesian inference. Fully embracing HPC technologies can unlock performance levels far beyond what current dataflow frameworks typically provide. In the following sections, we delve deeper into these challenges and present a roadmap for making MPI+X approaches more accessible, scalable, and impactful for the SC community.
@@ 2 HPC and SC: A History
In recent years, the evolution of computing hardware has entered a new phase, driven by the slowing of transistor scaling and the growing demand for high-performance processing. This marks a turning point from the earlier era when Moore’s Law reliably predicted rapid gains in transistor density. In 1965, Gordon Moore, co-founder of Intel, extrapolated four years of data to predict an exponential increase in the number of transistors per integrated circuit (Moore, 1965). The ensuing decades bore him out with a doubling period of approximately 18 months. However, that pace has since slowed. By 2005, the rise in single-chip central processing unit (CPU) performance, powered by millions of transistors, began to face serious thermal and power limitations [Sutter, 2005]. In response, hardware designers shifted focus from increasing clock speeds to integrating multiple cores on a single chip, introducing the era of multicore processors. These architectures introduced multiprocessing execution, placing new demands on software to manage parallelism effectively. In parallel with the shift toward multicore CPUs, another revolution was quietly taking shape. At the end of the 20th century, NVIDIA introduced the first standalone graphics processing unit (GPU), the “GeForce 256”, designed to accelerate gaming graphics [NVIDIA, 1999]. What began as a specialized tool for rendering images quickly evolved into a powerhouse for general-purpose computing. With far greater parallelism and memory bandwidth than traditional CPUs, GPUs have become indispensable accelerators in modern HPC, at the expense of less memory per processor. Today, GPUs play a central role not only in gaming but also in high-performance applications, including deep learning, scientific simulation, and large-scale statistical inference. Modern HPC systems capitalize on this capability, combining GPUs with multicore CPUs to solve problems that exceed the limits of conventional computing. These platforms are specifically designed to address computational challenges that require vast processing power, high memory throughput, and low-latency communication. Supercomputers, the summit of HPC infrastructure, integrate thousands of CPUs and GPUs connected by ultra-fast interconnects such as InfiniBand [Grun, 2010] and Slingshot [De Sensi et al., 2020]. This tightly coupled architecture enables rapid coordination across processing units, allowing for the efficient execution of large-scale simulations and data-intensive workloads. Beyond traditional processors, modern HPC systems increasingly incorporate specialized accelerators, such as Tensor Processing Units [TPUs, Jouppi et al., 2017] and Field-Programmable Gate Arrays [FPGAs, Putnam et al., 2014], to optimize specific tasks, including deep learning inference and cryptographic operations.
@@
The modern era of HPC began with the 1976 debut of the Cray‑1, which delivered roughly 133 MFLOPS. Progress accelerated quickly, producing more HPC machines. The Cray X-MP (1982) reached 800 MFLOPS, the Intel iPSC Hypercube (1985) surpassed 1.6 GFLOPS, the Cray Y-MP (1988) hit 2.67 GFLOPS, and the CM-5 (1991) achieved 131 GFLOPS [Dongarra et al., 1990]. These advances fueled breakthroughs across science, defense, and industry, introducing HPC as a foundational technology for large-scale computation. Over the past three decades, technologically leading nations have engaged in an ongoing race to develop and possess the world’s fastest supercomputers, as represented in the biannual TOP500 rankings [TOP500, 2024]. These machines have redefined what is computationally possible, powering breakthroughs in many applications, including physics, climate modeling, AI, and engineering. Their rise has been fueled by rapid innovations in CPU, GPU, and system architecture, often driven by advancements in the industry. Figure 2 shows this trajectory from 1993-2024, highlighting the top-performing supercomputers at each point in time. Peak performance, expressed on a logarithmic scale in FLOPS, highlights the exponential growth trend that has characterized HPC. From gigaflops in the early 1990s to exaflops today, the curve reveals a consistent doubling of computational power roughly every 1.5 to 2 years, mirroring Moore’s Law and the scaling of parallelism. Some systems, such as ASCI Red, BlueGene/L, Summit, Fugaku, and most recently Frontier and El Capitan, mark significant inflection points in this evolution. If this trend can be sustained, zettaflop-scale systems are expected to emerge as early as the 2030s, introducing a new era of extreme-scale computing for many scientific fields.
[…]
@@
figureEvolution of Supercomputer Peak Performance Over Time: From CM-5 to El Capitan.
@@
SC also has a long history dating back to the mid-20th century. One of the earliest documented statistical packages, the Bio-Medical Data Package (BMDP), was developed in the 1960s at the UCLA Health Computing Facility for IBM mainframes [Dixon, 1983]. Designed for parametric and nonparametric statistical analyses, BMDP marked the first wave of statistical computing. The second wave emerged in the late 1960s and early 1970s with the release of SPSS (1968), Genstat (1970), and SAS (1971) [Nie et al., 1970]. These packages introduced powerful tools such as regression, ANOVA, and cluster analysis to mainframe computing, revolutionizing statistical analysis in research, and were written in FORTRAN. The third statistical computing wave began with the development of the S language in the 1970s, led by John Chambers and his team at Bell Labs [Chambers, 2008]. Originally developed as a general-purpose statistical tool for academic users, S was later acquired by Insightful Corporation and rebranded as S-PLUS, which emerged as a leading commercial statistical software in the 1990s. Around the same time, Robert Gentleman and Ross Ihaka introduced R [Ihaka and Gentleman, 1996], establishing a new standard for open, extensible, and efficient statistical computing. Their work, published in the Journal of Computational and Graphical Statistics, laid the foundation for modern statistical software and established key principles for efficient statistical computation that remain relevant today. The late 1990s and early 2000s saw the convergence of this foundation with advances in parallel architectures and algorithms, setting the stage for a new chapter. This paper introduces the fourth wave in SC, i.e., HPSC. Unlike previous waves, HPSC is not defined by a specific language or software, but by a broader shift toward scalable and parallel approaches that leverage the power of modern HPC systems.
@@
Today, HPC includes a broad spectrum of technologies, most of which involve some form of parallel computing. Two dominant tracks have emerged: one rooted in the scientific community, the other in industry. In science, HPC is structured by national supercomputing centers focused on simulation science, typically leveraging the MPI+X programming model [Bernholdt et al., 2020, Hori et al., 2021]. Interestingly, Hori et al. [2021] uses R for portions of their data analysis and graphics. In contrast, the business world has adopted large-scale data centers, driven by cloud computing, search engines, and e-commerce, which favor dataflow-based parallelism [Kamburugamuve et al., 2018]. These divergences reflect fundamental differences in application needs, as well as a degree of cultural and technical isolation between the two communities. At first glance, one might assume the dataflow model is better suited to statistical computing due to its focus on data processing. However, the reality is more complex and in many cases, the opposite holds true. A careful comparison of three widely used algorithms in computational statistics [^1] Referred to as machine learning algorithms in the source, reflecting its computer science perspective. shows that MPI-based implementations consistently outperform their dataflow counterparts, often by significant margins [Kamburugamuve et al., 2018]. The MPI model, grounded in mathematical structures such as matrix algebra, aligns naturally with the computational patterns of statistical analysis. While dataflow systems offer strengths in scalability and data management, especially in long-running services, they are optimized for use cases distinct from the demands of HPSC.
@@ 3.1 Climate Science
HPSC has significantly advanced climate science by enabling scalable inference, prediction, and uncertainty quantification for massive spatio-temporal environmental datasets. In the literature, a notable contribution is the ExaGeoStat framework [Abdulah et al., 2018a, b, 2019], which introduced tile-based, architecture-agnostic solvers for dense and approximate Gaussian process (GP) likelihoods, supporting exact and approximate maximum likelihood estimation (MLE) for spatio-temporal data modeling on CPU-GPU clusters. Tile algorithms partition matrices into cache-sized tiles, enabling efficient multicore/GPU parallelism [Gustavson, 2001, Buttari et al., 2009].
@@
A key innovation in ExaGeoStat is the integration of tile low-rank (TLR) approximations, where each matrix tile is individually compressed in low-rank [Abdulah et al., 2018b], with mixed-precision techniques, where each tile is computed at a precision level (FP64/FP32/FP16) chosen adaptively based on matrix norms or application-specific features [Abdulah et al., 2019, Cao et al., 2022, 2023]. This approach reduces memory footprint and computational cost while preserving likelihood fidelity [Cao et al., 2022].
@@
Later extensions to ExaGeoStat pushed beyond Gaussian assumptions [Mondal et al., 2022, 2023], added hybrid-precision solvers with dynamic scheduling via PaRSEC [Abdulah et al., 2021], and enabled containerized deployments across diverse HPC systems [Abdulah et al., 2024b]. These innovations enabled large-scale inference on existing supercomputers with notable speedups in modeling evapotranspiration fields. To improve usability, ExaGeoStat was wrapped as the R package ExaGeoStatR [Abdulah et al., 2023], bridging high-performance modeling and the R computing environment. Within this ecosystem, Salvaña et al. [2021, 2022] developed massively parallel MLE solvers for multivariate and spatio-temporal GPs, leveraging distributed dense linear algebra and global parallel particle swarm optimization [Schutte et al., 2004] to support exascale kriging for climate and pollution data. Geng et al. [2023, 2025] further enhanced performance by implementing a GPU-accelerated modified Bessel function, which is integrated into ExaGeoStat for the efficient evaluation of the Matérn covariance function. Other ExaGeoStat features include convolutional neural networks-based partitioning of nonstationary spatial domains [Nag et al., 2025], diagnostics for predictive efficiency in low-rank models [Hong et al., 2021], and large-scale simulation capabilities showcased in three international competitions [Huang et al., 2021, Abdulah et al., 2022, Hong et al., 2023], which benchmarked performance across a range of covariance structures and approximations. ExaGeoStat’s contribution to HPSC in climate science was recognized when Abdulah et al. [2024a] received the ACM Gordon Bell Prize for Climate Modeling.
@@
Beyond ExaGeoStat, HPSC-enabled tools have advanced climate modeling through GPU-accelerated profile likelihoods for dense spatial covariance matrices [Xu and Brown, 2024], decentralized low-rank inference [Shi et al., 2025], sparsity‑discovering kernels [Noack et al., 2023], parallel particle-based Bayesian estimation [Katzfuss and Hammerling, 2017], and parallel Markov Chain Monte Carlo (MCMC) techniques [Solonen et al., 2012]. Moreover, Abdul Fattah et al. [2025] introduced sTiles, a GPU-accelerated framework for sparse Cholesky factorization of structured matrices, with a particular focus on block arrowhead types commonly arising in Bayesian inference and scientific computing.
@@
A GPU-powered neural network has also demonstrated significant potential for fast estimation of spatial covariance parameters [Gerber and Nychka, 2021]. These methods address scale and communication bottlenecks while supporting robust, interpretable inference [Mann, 2020]. Additional contributions include an out-of-core GPU-based geostatistical interpolation [Allombert et al., 2014], a parallel kriging technique based on the k-d tree method [Wei et al., 2015], and GPU-accelerated likelihood estimation software in R. Efficiency improvements in accelerating Cholesky decomposition using a mixed-precision approach [Cao et al., 2023], GPU implementation of Vecchia approximations [Pan et al., 2024, James and Guinness, 2024, Pan et al., 2025b, a], and parallel approaches to inference of spatial extremes [Castruccio et al., 2016] further illustrate how HPSC reshapes both methodology and practice in climate science.

### v3

@@ Abstract
We recognize the emergence of a statistical computing community focused on working with large computing platforms and producing software and applications that exemplify high-performance statistical computing (HPSC). The statistical computing (SC) community develops software that is widely used across disciplines. However, it remains largely absent from the high-performance computing (HPC) landscape, particularly on platforms such as those featured on the
www.top500.org or Green500 lists. Many disciplines already participate in HPC, mostly centered around simulation science, although data-focused efforts under the artificial intelligence (AI) label are gaining popularity. Bridging this gap requires both community adaptation and technical innovation to align statistical methods with modern HPC technologies. We can accelerate progress in fast and scalable statistical applications by building strong connections between the SC and HPC communities. We present a brief history of SC, a vision for how its strengths can contribute to statistical science in the HPC environment (such as HPSC), the challenges that remain, and the opportunities currently available, culminating in a possible roadmap toward a thriving HPSC community.
keywords:
GPUs, high-performance computing, mixed-precision computing, parallel statistical algorithms, statistical computing
@@ 1 Introduction
Statistical computing (SC) is a foundational discipline that aims to merge statistical theory with computational techniques to turn data into actionable insights [140]. At its core, it involves designing and implementing novel algorithms, simulations, and models to solve complex problems across various domains, including climate science, economics, and machine learning. Unlike general-purpose computing, statistical computing is grounded in data analysis, probabilistic reasoning, and statistical inference, drawing heavily on probability theory, statistics, and numerical methods [89]. As data continues to scale in both size and complexity, statistical computing has become indispensable, not only for managing large datasets but also for enabling predictive modeling and informed decision-making in real-time. Programming languages such as R [123], Julia [35], and Python [254] have played a critical role in this evolution, empowering researchers to rapidly post-process, analyze, and visualize data. Together, these tools have made SC a central engine of modern data science.
@@
To date, a lot more work within the SC community has been invested in adopting dataflow technologies, i.e., frameworks such as Apache Spark [101], Dask [212], or TensorFlow [1], that represent computations as directed acyclic graphs of operations on data than in exploring message passing interface (MPI)+X approaches, i.e., a hybrid parallel programming paradigm that combines MPI [92] for distributed-memory communication with an additional model “X” (such as OpenMP [50] for multicore CPUs or CUDA for GPUs) for shared-memory or accelerator-level parallelism [117], despite the latter being better suited for many of the most computationally demanding tasks in the field. There are multiple factors behind this trend. First, dataflow architectures are often perceived as more accessible, thanks to their relatively simple parallel execution models, in contrast to MPI-based systems, which require explicit message passing and distributed memory management. Such frameworks abstract away much of this complexity. This helps statisticians use them even if they do not fully understand parallel programming. Second, the need for interactive computing has strongly influenced choices. Tools like R and Python are deeply tied to exploratory, iterative analysis, where users develop models, visualize outputs, and debug interactively. Moreover, dataflow systems integrate more naturally into cloud-based and batch-processing environments, making them a compelling choice for both industry and large-scale analytics. Third, most statisticians and data scientists primarily work with multicore processors rather than large HPC clusters. This has led to a strong emphasis on shared-memory parallelism (e.g., OpenMP, threading in R and Python) rather than distributed memory paradigms, such as MPI.
While dataflow technologies offer some distributed computing capabilities, they are often employed in a manner that still aligns with a multicore approach rather than fully leveraging the power of supercomputing. Even when statisticians engage with distributed computing, it is often through tools such as Dask [212], Ray [181], or Spark, which offer a gentler learning curve and allow scaling across nodes without requiring low-level programming expertise. While these tools have clear advantages in usability and accessibility, we confirm that MPI+X remains significantly underutilized in SC. Supercomputing environments offer high efficiency, lower latency, and access to highly optimized numerical libraries. These advantages can substantially accelerate tasks such as large-scale simulations, Monte Carlo methods, and Bayesian inference. Fully embracing HPC technologies can unlock performance levels far beyond what current dataflow frameworks typically provide. In the following sections, we delve deeper into these challenges and present a roadmap for making MPI+X approaches more accessible, scalable, and impactful for the SC community.
@@ 2 HPC and SC: A History
In recent years, the evolution of computing hardware has entered a new phase, driven by the slowing of transistor scaling and the growing demand for high-performance processing. This marks a turning point from the earlier era when Moore’s Law reliably predicted rapid gains in transistor density. In 1965, Gordon Moore, co-founder of Intel, extrapolated four years of data to predict an exponential increase in the number of transistors per integrated circuit (Moore, 1965). The ensuing decades bore him out with a doubling period of approximately 18 months. However, that pace has since slowed. By 2005, the rise in single-chip central processing unit (CPU) performance, powered by millions of transistors, began to face serious thermal and power limitations [247]. In response, hardware designers shifted focus from increasing clock speeds to integrating multiple cores on a single chip, introducing the era of multicore processors. These architectures introduced multiprocessing execution, placing new demands on software to manage parallelism effectively. In parallel with the shift toward multicore CPUs, another revolution was quietly taking shape. At the end of the 20th century, NVIDIA introduced the first standalone graphics processing unit (GPU), the “GeForce 256”, designed to accelerate gaming graphics [191]. What began as a specialized tool for rendering images quickly evolved into a powerhouse for general-purpose computing. With far greater parallelism and memory bandwidth than traditional CPUs, GPUs have become indispensable accelerators in modern HPC, at the expense of less memory per processor. Today, GPUs play a central role not only in gaming but also in high-performance applications, including deep learning, scientific simulation, and large-scale statistical inference. Modern HPC systems capitalize on this capability, combining GPUs with multicore CPUs to solve problems that exceed the limits of conventional computing. These platforms are specifically designed to address computational challenges that require vast processing power, high memory throughput, and low-latency communication. Supercomputers, the summit of HPC infrastructure, integrate thousands of CPUs and GPUs connected by ultra-fast interconnects such as InfiniBand [93] and Slingshot [65]. This tightly coupled architecture enables rapid coordination across processing units, allowing for the efficient execution of large-scale simulations and data-intensive workloads. Beyond traditional processors, modern HPC systems increasingly incorporate specialized accelerators, such as Tensor Processing Units [TPUs, 135] and Field-Programmable Gate Arrays [FPGAs, 208], to optimize specific tasks, including deep learning inference and cryptographic operations.
@@
The modern era of HPC began with the 1976 debut of the Cray‑1, which delivered roughly 133 MFLOPS. Progress accelerated quickly, producing more HPC machines. The Cray X-MP (1982) reached 800 MFLOPS, the Intel iPSC Hypercube (1985) surpassed 1.6 GFLOPS, the Cray Y-MP (1988) hit 2.67 GFLOPS, and the CM-5 (1991) achieved 131 GFLOPS [74]. These advances fueled breakthroughs across science, defense, and industry, introducing HPC as a foundational technology for large-scale computation. Over the past three decades, technologically leading nations have engaged in an ongoing race to develop and possess the world’s fastest supercomputers, as represented in the biannual TOP500 rankings [253]. These machines have redefined what is computationally possible, powering breakthroughs in many applications, including physics, climate modeling, AI, and engineering. Their rise has been fueled by rapid innovations in CPU, GPU, and system architecture, often driven by advancements in the industry. Figure 1 shows this trajectory from 1993-2024, highlighting the top-performing supercomputers at each point in time. Peak performance, expressed on a logarithmic scale in FLOPS, highlights the exponential growth trend that has characterized HPC. From gigaflops in the early 1990s to exaflops today, the curve reveals a consistent doubling of computational power roughly every 1.5 to 2 years, mirroring Moore’s Law and the scaling of parallelism. Some systems, such as ASCI Red, BlueGene/L, Summit, Fugaku, and most recently Frontier and El Capitan, mark significant inflection points in this evolution. If this trend can be sustained, zettaflop-scale systems are expected to emerge as early as the 2030s, introducing a new era of extreme-scale computing for many scientific fields.
@@
Figure 1: Evolution of Supercomputer Peak Performance Over Time: From CM-5 to El Capitan.
@@
SC also has a long history dating back to the mid-20th century. One of the earliest documented statistical packages, the Bio-Medical Data Package (BMDP), was developed in the 1960s at the UCLA Health Computing Facility for IBM mainframes [71]. Designed for parametric and nonparametric statistical analyses, BMDP marked the first wave of statistical computing. The second wave emerged in the late 1960s and early 1970s with the release of SPSS (1968), Genstat (1970), and SAS (1971) [189]. These packages introduced powerful tools such as regression, ANOVA, and cluster analysis to mainframe computing, revolutionizing statistical analysis in research, and were written in FORTRAN. The third statistical computing wave began with the development of the S language in the 1970s, led by John Chambers and his team at Bell Labs [49]. Originally developed as a general-purpose statistical tool for academic users, S was later acquired by Insightful Corporation and rebranded as S-PLUS, which emerged as a leading commercial statistical software in the 1990s. Around the same time, Robert Gentleman and Ross Ihaka introduced R [123], establishing a new standard for open, extensible, and efficient statistical computing. Their work, published in the Journal of Computational and Graphical Statistics, laid the foundation for modern statistical software and established key principles for efficient statistical computation that remain relevant today. The late 1990s and early 2000s saw the convergence of this foundation with advances in parallel architectures and algorithms, setting the stage for a new chapter. This paper introduces the fourth wave in SC, i.e., HPSC. Unlike previous waves, HPSC is not defined by a specific language or software, but by a broader shift toward scalable and parallel approaches that leverage the power of modern HPC systems.
@@
Today, HPC includes a broad spectrum of technologies, most of which involve some form of parallel computing. Two dominant tracks have emerged: one rooted in the scientific community, the other in industry. In science, HPC is structured by national supercomputing centers focused on simulation science, typically leveraging the MPI+X programming model [34, 117]. Interestingly, [117] uses R for portions of their data analysis and graphics. In contrast, the business world has adopted large-scale data centers, driven by cloud computing, search engines, and e-commerce, which favor dataflow-based parallelism [136]. These divergences reflect fundamental differences in application needs, as well as a degree of cultural and technical isolation between the two communities. At first glance, one might assume the dataflow model is better suited to statistical computing due to its focus on data processing. However, the reality is more complex and in many cases, the opposite holds true. A careful comparison of three widely used algorithms in computational statistics [^1] Referred to as machine learning algorithms in the source, reflecting its computer science perspective. shows that MPI-based implementations consistently outperform their dataflow counterparts, often by significant margins [136]. The MPI model, grounded in mathematical structures such as matrix algebra, aligns naturally with the computational patterns of statistical analysis. While dataflow systems offer strengths in scalability and data management, especially in long-running services, they are optimized for use cases distinct from the demands of HPSC.
@@ 3.1 Climate Science
HPSC has significantly advanced climate science by enabling scalable inference, prediction, and uncertainty quantification for massive spatio-temporal environmental datasets. In the literature, a notable contribution is the ExaGeoStat framework [9, 10, 11], which introduced tile-based, architecture-agnostic solvers for dense and approximate Gaussian process (GP) likelihoods, supporting exact and approximate maximum likelihood estimation (MLE) for spatio-temporal data modeling on CPU-GPU clusters. Tile algorithms partition matrices into cache-sized tiles, enabling efficient multicore/GPU parallelism [96, 41].
@@
A key innovation in ExaGeoStat is the integration of tile low-rank (TLR) approximations, where each matrix tile is individually compressed in low-rank [10], with mixed-precision techniques, where each tile is computed at a precision level (FP64/FP32/FP16) chosen adaptively based on matrix norms or application-specific features [11, 44, 45]. This approach reduces memory footprint and computational cost while preserving likelihood fidelity [44].
@@
Later extensions to ExaGeoStat pushed beyond Gaussian assumptions [178, 179], added hybrid-precision solvers with dynamic scheduling via PaRSEC [6], and enabled containerized deployments across diverse HPC systems [7]. These innovations enabled large-scale inference on existing supercomputers with notable speedups in modeling evapotranspiration fields. To improve usability, ExaGeoStat was wrapped as the R package ExaGeoStatR [8], bridging high-performance modeling and the R computing environment. Within this ecosystem, [222, 223] developed massively parallel MLE solvers for multivariate and spatio-temporal GPs, leveraging distributed dense linear algebra and global parallel particle swarm optimization [229] to support exascale kriging for climate and pollution data. [86, 87] further enhanced performance by implementing a GPU-accelerated modified Bessel function, which is integrated into ExaGeoStat for the efficient evaluation of the Matérn covariance function. Other ExaGeoStat features include convolutional neural networks-based partitioning of nonstationary spatial domains [186], diagnostics for predictive efficiency in low-rank models [115], and large-scale simulation capabilities showcased in three international competitions [122, 4, 116], which benchmarked performance across a range of covariance structures and approximations. ExaGeoStat’s contribution to HPSC in climate science was recognized when [5] received the ACM Gordon Bell Prize for Climate Modeling.
@@
Beyond ExaGeoStat, HPSC-enabled tools have advanced climate modeling through GPU-accelerated profile likelihoods for dense spatial covariance matrices [269], decentralized low-rank inference [232], sparsity‑discovering kernels [190], parallel particle-based Bayesian estimation [138], and parallel Markov Chain Monte Carlo (MCMC) techniques [236]. Moreover, [3] introduced sTiles, a GPU-accelerated framework for sparse Cholesky factorization of structured matrices, with a particular focus on block arrowhead types commonly arising in Bayesian inference and scientific computing.
@@
A GPU-powered neural network has also demonstrated significant potential for fast estimation of spatial covariance parameters [88]. These methods address scale and communication bottlenecks while supporting robust, interpretable inference [169]. Additional contributions include an out-of-core GPU-based geostatistical interpolation [14], a parallel kriging technique based on the k-d tree method [261], and GPU-accelerated likelihood estimation software in R. Efficiency improvements in accelerating Cholesky decomposition using a mixed-precision approach [45], GPU implementation of Vecchia approximations [201, 130, 202, 200], and parallel approaches to inference of spatial extremes [48] further illustrate how HPSC reshapes both methodology and practice in climate science.